OpenStudy (anonymous):

Medal given to best answer!! Zen wants to fill a window box with soil. The length of the window box is 5 feet, width is 6 inches, and height is 9 inches. Each bag of soil contains 0.625 cubic foot of soil. How many bags of soil will Zen need to fill the window box completely? [1 foot = 12 inches]

OpenStudy (anonymous):

do the volume of the box in cubic feet: V=5*0.5*0.75 = 15/8 Volume of soil bag is Vbag=0.625=5/8 Hopefully, you can see the rest. Hint 3

OpenStudy (anonymous):

@nervegrinder?

OpenStudy (anonymous):

He needs 3 bags?